OpenStudy (mony01):

A model for the basal metabolism rate in kcal/h, of a young man is R(t)=85-0.18cos(pi(t)/12), where t is the time in hours measured from 5:00 am. What is the total basal metabolism of this man, over a 24-hour time period?

OpenStudy (mony01):

\[\int\limits_{0}^{24}R(t) dt\]

ganeshie8 (ganeshie8):

looks good. the phrase 'time is measured from 5:00am' is begginhog us to set bounds as 5->29 but u wil get the same answer so it should be okay

OpenStudy (mony01):

so the integral is \[\int\limits_{0}^{24}85-0.18\cos(\frac{ \pi(t) }{12 })\]

ganeshie8 (ganeshie8):

yes, but i prefer setting it up as : \( \int\limits_{5}^{29}85-0.18\cos(\frac{ \pi(t) }{12 }) dt \)

OpenStudy (mony01):

is it 5 to 24?

ganeshie8 (ganeshie8):

5 to 29

OpenStudy (mony01):

what would the answer be?

OpenStudy (mony01):

can you tell me if this is the set up when you take the anti-derivative 85t-0.18sin(pi(t)/12)^2/2 dt

ganeshie8 (ganeshie8):

\(\large \int\limits_{5}^{29}85-0.18\cos(\frac{ \pi(t) }{12 }) dt\) \(\large \int\limits_{5}^{29}85 dt -\int\limits_{5}^{29} 0.18\cos(\frac{ \pi(t) }{12 }) dt\) \(\large 85t \Big|_5^{29} -\frac{0.18\sin(\frac{ \pi(t) }{12 })}{\frac{\pi}{12}} \Big|_5^{29}\)

ganeshie8 (ganeshie8):

take the limits now